OpenStudy (anonymous):

Fiona has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $25.10. How many quarters and how many nickels are in the jar? A. 88 quarters; 62 nickels B. 80 quarters; 70 nickels C. 75 quarters; 75 nickels D. 62 quarters; 88 nickels

OpenStudy (anonymous):

help please!!

OpenStudy (ranga):

Assume x quarters and y nickels. x + y = 150 (1) 0.25x + 0.05y = 25.10 (2) Solve (1) and (2)

OpenStudy (anonymous):

ok

OpenStudy (ranga):

Do you know how to solve (1) and (2) by elimination?

OpenStudy (ranga):

To eliminate y, multiply first equation by -0.05 and add it to the second equation.

OpenStudy (anonymous):

okay im following you

OpenStudy (anonymous):

sorry i dont get it that well

OpenStudy (ranga):

x + y = 150 (1) 0.25x + 0.05y = 25.10 (2) To eliminate y, multiply (1) by -0.05 and add it to (2): -0.05x - 0.05y = -7.5 0.25x + 0.05y = 25.10 (2) add: 0.2x = 17.6 x = 17.6 / 0.2 = 88 put x = 88 in (1): 88 + y = 150 y = 150 - 88 = 62 There are 88 quarters and 62 nickels.